### v1

# The impact of gravitational waveform model systematics on the measurement of the Hubble Constant

Abstract
Matching gravitational-wave observations of binary neutron stars with theoretical model predictions reveals important information about the sources, such as the masses and the distance to the stars. The latter can be used to determine the Hubble Constant, the rate at which the Universe expands. One general problem of all astrophysical measurements is that theoretical models only approximate the real underlying physics, which can lead to systematic uncertainties introducing biases. However, the extent of this bias for the distance measurement due to uncertainties of gravitational waveform models is unknown. In this study, we analyze a synthetic population of 38 binary neutron star sources measured with Advanced LIGO and Advanced Virgo at design sensitivity. We employ a set of four different waveform models and estimate model-dependent systematic biases on the extraction of the Hubble Constant using the bright siren method. Our results indicate that systematic biases are below statistical uncertainties for the current generation of gravitational wave detectors.
I Introduction
The measurement of the local expansion rate of the Universe, also known as the Hubble Constant, $H_{0}$, is one of the most intriguing problems in cosmology. The knowledge of the cosmological expansion rate at different times, given through the Hubble-Lemaître parameter, is essential for understanding the history of the Universe and to infer its age and composition.
A big puzzle in cosmology is to piece together the measurements from the early and late Universe. While the SHOES collaboration Riess et al. (2019) using the cosmic distance ladder method constrained the local expansion rate to $H_{0}=74.03\pm 1.42\rm\ km\ s^{-1}Mpc^{-1}$, the Planck Collaboration Aghanim et al. (2020) estimated $H_{0}=67.4\pm 0.5\rm\ km\ s^{-1}Mpc^{-1}$ at 68% credibility from observations of the Cosmic Microwave Background (CMB) radiation. This leads to the well-known Hubble-Lemaître tension with a discrepancy at a $4.4$ sigma level Riess et al. (2019).
This discrepancy has spurred scientific efforts, with researchers proposing numerous independent studies to elucidate the underlying cause. The use of acoustic baryonic oscillations Macaulay et al. (2019), the gravitational lensing method Birrer et al. (2019), or calibrating the Tip of the Red Giant Branch (TRGB) to Type Ia supernovae Freedman et al. (2019) exemplify prominent attempts to resolve this enigma. Meanwhile, the possibility of a need for new physics is actively debated, suggesting that a new cosmological model may be necessary to reconcile the distinct measurements Berbig et al. (2020); Blinov et al. (2022); Mandal et al. (2023); Poulin et al. (2023).
Using gravitational-wave (GW) observations as a probe to infer the cosmic expansion rate was first proposed by Schutz in 1986 Schutz (1986). The GW emission of compact binary coalescences involving black holes or neutron stars can serve as cosmic distance indicators or so-called standard sirens and provide a measurement of the luminosity distance of the source. The landmark detection of GW170817 Abbott et al. (2017a, 2019), a binary neutron star (BNS) merger, by the LIGO and Virgo detectors Aasi et al. (2015a); Acernese et al. (2015a) along with its associated electromagnetic (EM) counterparts, GRB170817A and AT2017gfo (Abbott et al. (2017b) and refs. therein), marked a pivotal milestone in the field of cosmology. This event enabled the first direct measurement of the Hubble constant using a GW standard siren Abbott et al. (2017c). Although current constraints from GW observations are not yet able to resolve the Hubble tension, future observations could place further constraints on this cosmological parameter Chen et al. (2018); Feeney et al. (2019); Farr et al. (2019a); Feeney et al. (2021). Especially multi-messenger observations hold significant promise for improving the accuracy of Hubble constant determinations. This advantage arises from being able to combine the GW measurement of the distance to the source with an electromagnetic measurement of its redshift Abbott et al. (2017c); Guidorzi et al. (2017); Hotokezaka et al. (2019); Wang and Giannios (2021); Dietrich et al. (2020); Bulla et al. (2022); Palmese et al. (2024).
While the standard siren methodology offers the advantage of not requiring re-calibration, as is common practice in the distance ladder method, there are different sources of systematic uncertainties that must be addressed. Previous studies such as Ref. Huang et al. (2022) investigated the impact of instrumental calibration uncertainties on Hubble constant measurements. They found that for single BNS events with a signal-to-noise ratio (SNR) of 50 for a network consisting of Advanced LIGO and Advanced Virgo operating at design sensitivity, the introduced bias is smaller than statistical uncertainties, but could accumulate as more events will be detected. Ref. Turski et al. (2023) examined the influence of galaxy redshift uncertainties using different galaxy redshift uncertainty models, with particular emphasis on photometric redshifts, which are known to exhibit substantial errors. Their findings indicate that while there may exist a potential bias, it remains smaller than statistical uncertainties. Furthermore, systematic uncertainties arising from the viewing angle of BNS mergers and EM selection effects can be major challenges Chen (2020); Chen et al. (2024); Mancarella et al. (2024). Building on Dhani et al. (2024), this work explores the systematic biases introduced by binary black hole (BBH) waveform models. The analysis reveals that these biases become more pronounced with increasing detector-frame total mass, binary asymmetry, and spin-precession effects. Notably, the study demonstrates on the basis of three high SNR events (golden events) that current waveform models lead to biased measurement of the Hubble constant, even for current detectors.
In contrast to Ref. Dhani et al. (2024), our study delves into the issue of systematic uncertainties originating from BNS waveform models potentially influencing the determination of the distance to the source and, hence, the Hubble constant. Until now, this question remains, up to the best of our knowledge, unanswered. We analyze a synthetic population of 38 BNS sources, previously published in Ref. Kunert et al. (2022), with a set of four different waveform models and study model-dependent systematic errors. We employ the most recent Planck CMB measurement Aghanim et al. (2020) as a benchmark and investigate whether estimating $H_{0}$ using different models than the one used to generate the data introduces a systematic bias. As our results will demonstrate, GW model dependent biases have a minimal influence on the inferred value of the Hubble constant, for current detector networks.
This paper is organized as follows. Section II provides details on modeling GW standard sirens, the simulated BNS source population, and on the Bayesian framework employed to extract estimates of $H_{0}$. Our results of combined $H_{0}$-estimates are presented in section III and section IV summarizes the key takeaways.
II Methodology
The standard siren method utilizes the detection of GW signals and their correlation with EM observations of the same astrophysical events to infer the Hubble constant. When a source’s redshift can be measured directly or through an association with its host galaxy, it is denoted as a bright siren. When an EM counterpart is absent, statistical methods can be used to ascertain possible redshifts of the source. Notable examples of these techniques include the galaxy catalog method Del Pozzo (2012); Chen et al. (2018); Fishbach et al. (2019); Gray et al. (2020); Abbott et al. (2023); Gray et al. (2022, 2023); Mastrogiovanni et al. (2023); Borghi et al. (2024), the cross correlation method Oguri (2016); Mukherjee et al. (2021); Bera et al. (2020), and the spectral siren technique Farr et al. (2019b); Mastrogiovanni et al. (2021); Mukherjee (2022); Ezquiaga and Holz (2022).
In our study, we will employ the bright siren technique to infer a joint estimate of the local expansion rate of the Universe, assuming the standard Lambda cold dark matter ($\Lambda$CDM) cosmology. In accordance, we utilize the latest Planck CMB measurement Aghanim et al. (2020) as a reference value for our Bayesian analysis.
II.1 Modelling standard sirens
Gravitational waveform models describe the amplitude $\tilde{A}$ and the phase $\phi$ of standard sirens emitted from coalescing compact objects like neutron stars and black holes. The evolution of a gravitational waveform $\tilde{h}(f)$ over frequency can be written as
$\tilde{h}(f)=\tilde{A}(f)\exp(-i\phi(f)),$ (1)
in which both, the amplitude as well as the phase carry information about the intrinsic source properties. GW detectors measure a linear combination of two polarizations of the GW signal. At leading order, these can be written as Gray et al. (2020)
$\tilde{h}_{+}(f)\propto\frac{\mathcal{M}_{z}^{5/6}}{D_{L}}\frac{1+\cos^{2}(\iota)}{2}f^{-7/6}\exp({i\phi(\mathcal{M}_{z},f)}),$ (2)
$\tilde{h}_{\times}(f)\propto\frac{\mathcal{M}_{z}^{5/6}}{D_{L}}\cos{\iota}f^{-7/6}\exp({i\phi(\mathcal{M}_{z},f)+i\pi/2}),$ (3)
where $\mathcal{M}_{z}\equiv\mathcal{M}(1+z)$ denotes the detector-frame (or “red-shifted”) chirp mass, which is better constrained by the signal’s phase as compared its amplitude. It can be seen that the amplitude of a GW signal decreases in proportion to the luminosity distance $D_{L}$ of its source. This inverse relationship is a fundamental property of GWs and plays a crucial role in their detection and interpretation. On the other hand, gravitational waveform models including only the quadrupole mode show a limited capability to accurately infer the luminosity distance to an astrophysical source due to its interdependence with the inclination angle $\iota$. This degeneracy fundamentally limits our ability to measure both parameters so that the dominant uncertainty in the signal’s amplitude arises from the uncertainties on the luminosity distance and the inclination angle. The inclusion of higher-order modes can be a way to break this degeneracy, as shown in Ref. Calderón Bustillo et al. (2021).
II.2 Population of synthetic BNS sources
In this study, we use the large set of mock data obtained in Ref. Kunert et al. (2022) and exploit the inferred posterior on the luminosity distance to study the impact of model uncertainties on the estimation of the Hubble Constant. Below, we recap the key elements of the injection study.
The population of $38$ synthetic binary neutron star sources were injected with stationary Gaussian noise into a network of interferometers assuming Advanced LIGO Aasi et al. (2015b) and Advanced Virgo Acernese et al. (2015b) design sensitivity. In light of current observations suggesting a uniform distribution for neutron star masses in gravitational-wave binaries Landry and Read (2021), we adopt a uniform prior sampled within $m_{1,2}\in[1,2]~{}M_{\odot}$. All sources were injected uniformly in co-moving volume with an optimal network SNR, $\rho$, ranging within $\rho\in[7,100]$ and lie within a maximum or threshold luminosity distance of $D^{\rm thr}_{L}=100$ Mpc. Hence, the population lies within a volume in which cosmology can be well modelled under the linear Hubble relation (for redshifts $z\ll 1$) given as
$D_{L}(z)\approx\frac{cz}{H_{0}},$ (4)
in which $c$ is the speed of light. For all BNS signals in our population, we employed IMRPhenomD_NRTidalv2 (PhenDNRTv2) Dietrich et al. (2019a) as our injection model. We performed parameter estimation runs using parallel bilby Smith et al. (2020) for all sources in the population with four different models, which were TaylorF2 (TF2) Sathyaprakash and Dhurandhar (1991); Blanchet et al. (1995); Damour et al. (2001); Blanchet et al. (2004, 2005); Mishra et al. (2016); Mikoczi et al. (2005); Arun et al. (2009); Bohé et al. (2015); Mishra et al. (2016); Damour and Nagar (2010); Vines et al. (2011); Bini et al. (2012); Damour et al. (2012), IMRPhenomD_NRTidal (PhenDNRT) Husa et al. (2016); Khan et al. (2016); Taracchini et al. (2014); Dietrich et al. (2017, 2019b); Nagar et al. (2018); Bernuzzi et al. (2015), IMRPhenomD_NRTidalv2 (PhenDNRTv2), and SEOBNRv4_ROM_NRTidalv2 (SEOBNRTv2) Bohé et al. (2017a, b); Dietrich et al. (2019a) resulting in a total of 152 simulations. For further details on our injection setup, we refer the interested reader to Ref. Kunert et al. (2022).
II.3 Bayesian framework
In this section, we outline our method for estimating the Hubble constant by utilizing the posterior distributions on luminosity distance obtained for the discussed BNS population. In order to compute the posterior probability distribution of the Hubble constant for a set of GW events $\left\{x_{\rm{GW}}\right\}$, we employ Bayes’ theorem
$p(H_{0}|\left\{x_{\rm{GW}}\right\})\propto p(H_{0})p(\left\{x_{\rm{GW}}\right\}|H_{0})$ (5)
where $p(H_{0})$ is the prior on $H_{0}$ and $p(\left\{x_{\rm{GW}}\right\}|H_{0})$ is the likelihood of the measured GW data given a certain value of $H_{0}$.
Our analysis follows the bright siren approach, which utilizes the joint information from GW detections $\left\{x_{\rm{GW}}\right\}$ and electromagnetic observations providing information on the redshift $\left\{\hat{z}\right\}$ Abbott et al. (2017c). Consequently, we determine the joint posterior, $p(H_{0}|\left\{x_{\rm GW},\hat{z}\right\})$, for $N_{\rm obs}$ events in our population as follows
[…]
where $p(D_{L})$ is the prior on the luminosity distance and $p(\vec{\lambda})$ relates to priors for other GW parameters which are denoted as $\vec{\lambda}$. Analogous to the approach used in Ref.Abbott et al. (2017c), $N_{s}(H_{0})$ incorporates the selection effect imposed by the finite sensitivity of GW detectors and is
[…]
integrated over the prior probability distributions of the parameters $D_{L},\vec{\lambda}$ and over “detectable” GW data sets. This factor is constructed on the assumption that the choice of whether or not an event is included in the analysis is a property purely of the observed data, $x_{\rm GW}$. In this analysis (as in Abbott et al. (2017c)) we assume that selection effects are dominated by the GW data, i.e., there are no selection effects for the EM counterpart. Additionally, we assume that GW detection is based on the observed SNR of the source [^1] In practice, we impose selection on the true, rather than the observed, luminosity distance of the source. This is the approach used widely in the literature. Selection on source parameters rather than observed data should really be treated differently Gair et al. (2023), but in practice this approximation does not change the results.. At the distances to which BNS events can be observed with current GW detectors, such a selection becomes a threshold in the luminosity distance of the source, which is what the GW detectors measure. As a consequence, the selection function has no dependence on the unknown cosmological parameters and so can be set to unity.
Since our dataset lacks direct redshift measurements for each BNS merger event, we need to model this information. We consider all EM observations as securely detected and, hence, neglect a correction for EM selection effects. In the absence of actual electromagnetic observational data for redshifts, we simulate a redshift measurement $\hat{z}$ for the $i$-th BNS merger event using a Gaussian probability density function
$p(\hat{z}_{i}|D_{L},H_{0})=\frac{1}{\sqrt{2\pi}\sigma_{z}}\exp\Bigl{[}-\frac{1}{2}\Bigl{(}\frac{\hat{z}-z}{\sigma_{z}}\Bigr{)}^{2}\Bigr{]}.$ (8)
This allows us to associate each BNS source with a simulated measured redshift sampled from a Gaussian distribution centered on its true, injected redshift value. We use the same Gaussian likelihood in Eq. (6) to obtain our Hubble constant posteriors. We examine the dependence of our analysis on the choice of the fractional error term, $C$, which contributes to the overall redshift uncertainty, $\sigma_{z}$, defined by the linear relationship $\sigma_{z}=z\cdot C$. This aims to quantify the sensitivity of our results to variations in the assumed fractional error value and, hence, in the precision of the EM measurement of the redshift. To infer joint constraints on $H_{0}$, we convert the measured redshift values to luminosity distances using Eq. (4).
III Results
In accordance with the methodology detailed in section II.3, we obtained combined $H_{0}$-estimates for each GW model. As described, we aim to investigate the sensitivity of our results to variations in the fractional redshift error value $C$, which we explore for $C=0.5\%$ and $3\%$.
First of all, we study how our estimate on $H_{0}$ changes with the number of successively combined GW signals for each BNS event for $C=0.03$, i.e., a 3% fractional error in the modeling of the observed redshift. The result is shown in Fig. 1 and shows that all models recover the Planck injection at $H_{0}=67.4~{}\rm kms^{-1}Mpc^{-1}$ within their uncertainty bands at 95% credible interval. Consistent with theoretical expectations, incorporating more GW events gradually decreases the uncertainties associated with the $H_{0}$ estimates. Among 38 combined GW events, PhenDNRTv2 (the “true” model) and PhenDNRT align best with the injected value, while SEOBNRTv2 and TF2 show slightly larger deviations from the injection. From Fig. 1, we find that systematic biases in the extraction of the Hubble constant are below statistical uncertainties.
Figure 1: $H_{0}$-results versus the number of successively combined GW events for the GW models PhenDNRT (yellow), TF2 (purple), SEOBNRTv2 (orange), and the injection model PhenDNRTv2 (blue). Constraints on $H_{0}$ are reported at 95% credible interval. The injection value of the Planck measurement Aghanim et al. (2020) is shown as black dashed line. For this result, we modeled a $3$% uncertainty on the true, injected redshift as described in Sec. II.3.
Secondly, we show the results for a 0.5% fractional error in Fig. 2, i.e., assuming more precise EM measurements of the observed redshift. We find that our injection model, PhenDNRTv2, recovers the injection value of the Planck measurement within 95% credible interval. Likewise, analyses employing the GW models PhenDNRT and TF2 recover the injection signal when considering the combined data from all BNS events. The only exception is the model SEOBNRTv2 which does not recover the injection when data from roughly 30 BNS events have been combined. Moreover, we find that there is a slight overestimation for all models when comparing with the injected Planck measurement, most likely due to statistical fluctuations.
At this point, it is crucial to acknowledge that our recovered $H_{0}$ values presented in Fig. 2 depend upon the specific random realization of observed redshifts generated through Eq. (8). This implies that any variation in the random realization of redshift measurements for each event would lead to a corresponding change in the inferred $H_{0}$ posterior. While we present a single realization in Fig. 2, we investigated the dependence of our results with regard to different random realizations of the redshift measurement. Notably, while maintaining a fixed fractional error of 0.5%, we observed that the final, combined $H_{0}$ estimates exhibit a tendency to converge closer to the injected value under certain realizations (not shown in Fig. 2).
In order to minimize the impact of stochastic redshift realizations and of specific BNS event orderings in our sample, we implement a double randomization procedure [^2] Despite this procedure, our GW posteriors are not randomized, which can lead to slight offsets or biases as we can see in our study.. This involves shuffling the BNS event order for 100 times and drawing 100 random redshift realizations. The result is shown in Fig. 3, for which we modeled a 0.5% uncertainty on the true, injected redshift. Our result depicts a convergence towards the injected Planck value for all GW models. Consistent with prior findings, it also reveals a progressive reduction in uncertainties and that all GW models recover the injection.
Figure 2: $H_{0}$-results versus the number of successively combined GW events for the GW models PhenDNRT (yellow), TF2 (purple), SEOBNRTv2 (orange), and the injection model PhenDNRTv2 (blue). Constraints on $H_{0}$ are reported at 95% credible interval. The injection value of the Planck measurement Aghanim et al. (2020) is shown as black dashed line. For this result, we modeled a $0.5$% uncertainty on the true, injected redshift as described in Sec. II.3.
From Figs. 2 and 3, we find that PhenDNRT tends to yield smaller $H_{0}$-values as compared to PhenDNRTv2. This could result from the fact that the PhenDNRTv2 includes an amplitude tidal correction term Dietrich et al. (2019a), while this is not the case for PhenDNRT. Due to this correction, the amplitude for PhenDNRT waveforms will be larger as compared to PhenDNRTv2. In order to compensate, PhenDNRT needs to go to higher luminosity distances and, hence, will predict smaller values for the Hubble constant. Another distinctive feature is that the uncertainty is smallest for TF2 when all data has been taken into account. To investigate this observation, we computed the Fisher information matrix (FIM) for both TF2 and PhenDNRTv2 across all events using GWfast Iacovelli et al. (2022a, b). The FIM analysis revealed that 55% of the events exhibited a lower distance error for TF2 as compared to PhenDNRTv2. This finding suggests that TF2 can achieve a narrower posterior width or lower variance. However, we would still expect it to have a larger bias, which would eventually dominate the uncertainty once sufficiently many events have been combined. Moreover, we repeat the analysis employing the Hubble constant measurement reported by the SHOES collaboration Riess et al. (2019) as the injected value, $H_{0}=74.03\pm 1.42\rm\ km\ s^{-1}Mpc^{-1}$. Figure 3 shows that we obtain consistent results compared to the scenario where we utilize the Planck measurement. Consequently, our primary conclusions remain unchanged.
Overall, our results indicate that waveform model-dependent biases on the extraction of the Hubble constant are small for current detectors. Ascribing the subtle discrepancies observed in the combined $H_{0}$-estimates solely to specific model descriptions or assumptions presents a significant challenge. Firstly, GW models only possess a limited capability to accurately infer the distance and consequently, the Hubble constant, so we are dominated by statistical rather than systematic uncertainties. Secondly, our employed GW models exhibit interdependencies within specific regimes which hampers an unambiguous explanation attributable to individual models.
Figure 3: $H_{0}$-results versus the number of successively combined GW events for the GW models PhenDNRT (yellow), TF2 (purple), SEOBNRTv2 (orange), and the injection model PhenDNRTv2 (blue) for both, using the SHOES Riess et al. (2019) measurement of $H_{0}=74.03\pm 1.42\rm\ km\ s^{-1}Mpc^{-1}$ and the Planck measurement Aghanim et al. (2020) of $H_{0}=67.4\pm 0.5\rm\ km\ s^{-1}Mpc^{-1}$ as injected value. We employed the same credible interval and fractional error on the redshift uncertainty as in Fig. 2, but randomly shuffled the order of BNS events in our sample for 100 times. We employ 100 random realizations of redshift measurements. The injection value of the Planck measurement is shown as black dashed line, while the injection of the SHOES measurement Riess et al. (2019) is shown as black dotted line.
IV Conclusions
In this study, we investigated the impact of gravitational waveform systematics on the extraction of the Hubble constant. Employing the bright siren approach, we obtained combined estimates on the Hubble constant by using previous simulated GW data from a population containing 38 BNS sources published in Kunert et al. (2022) and by modeling a corresponding EM observation to provide information on the redshift. Overall, our analysis reveals that the uncertainty in the luminosity distance arising from GW models exerts a negligible influence on the determination of the Hubble constant for the current generation of GW detectors. However, our analysis is limited because our employed GW models are based on the quadrupolar mode and do not include any higher-order modes. While current systematic biases on the Hubble constant are below statistical uncertainties, this picture will change with upcoming third generation detectors like the Einstein Telescope (ET) Punturo et al. (2010); Maggiore et al. (2020); Branchesi et al. (2023) or the Cosmic Explorer (CE) Reitze et al. (2019); Evans et al. (2021). Projected BNS detection rates of $10^{4}$ BNS/year for ET Regimbau et al. (2012); Singh et al. (2022); Branchesi et al. (2023) or ${}10^{5}$ BNS/year for a network consisting of ET and CE Walker et al. (2024) imply diminishing statistical uncertainties, necessitating an increased focus on reducing systematic biases in the future.
Acknowledgements.
N.K., P.T.H.P., T.D. acknowledge support from the Daimler and Benz Foundation for the project “NUMANJI”. T.D. acknowledges support from the European Union (ERC, SMArt, 101076369). Views and opinions expressed are those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. P.T.H.P. is supported by the research program of the Netherlands Organisation for Scientific Research (NWO). We acknowledge usage of computer time on Lise/Emmy of the North German Supercomputing Alliance (HLRN) [project bbp00049], on HAWK at the High-Performance Computing Center Stuttgart (HLRS) [project GWanalysis 44189], and on SuperMUC NG of the Leibniz Supercomputing Centre (LRZ) [project pn29ba].